# A multiplexed control architecture for superconducting qubits with row-column addressing

### 06 Mar 2024 v1

@@ III Gate operations in the multiplexed control architecture
Here we turn to illustrate how single- and two-qubit gates are realized in the multiplexed control architecture. As shown in Figs. 1(c) and 2, gate operations, unlike that in the conventional independent control scheme, are implemented by applying simultaneous control pulses to all the shared row and column lines. When focusing on one particular qubit or coupler in the lattice, single- and two-qubit gate operations are realized by applying a unique pair of control pulses to a qubit or a coupler. In this section, we propose various gate schemes that are both compatible with the two-pulse configuration and the existing superconducting qubit technologies. For illustration purposes, we focus on the gates applied to transmon qubits Koch2007 , but in principle, it should also be feasible for other superconducting qubits, such as fluxonium qubits Manucharyan2009 .
@@ III.1 Single-qubit gates
Table 1: Single-qubit gate schemes with two control pulses. The schematic illustration of these schemes is shown in Fig. 3. Note that for frequency-tunable qubits, static DC flux biases (not listed here, but shown in Fig. 3) are generally needed for biasing the qubits at their idle points.
@@
Table 1 summarizes four single-qubit gate schemes using two control pulses, which are also schematically illustrated in Fig. 3. The schemes can be grouped into two categories, depending on whether the qubit is tunable. In the following, we shall give detailed descriptions of the four schemes.
Type-ST: This category is for single-qubit gates applied to frequency-tunable qubits. During single-qubit gate operations, a flux pulse and a microwave pulse are simultaneously delivered to the qubit through a shared Z line and a shared XY line, respectively, as shown in Fig. 3(a). Additionally, for frequency-tunable qubits, static DC flux biases are generally needed for biasing the qubits at their idle points. Following Ref. Johnson2010 , this flux offset can in principle be introduced by using $\Phi$-DAC, which functions as on-chip programmable magnetic memory, providing the needed static bias. Moreover, the $\Phi$-DAC can be efficiently programmed and addressed by examining the addressing circuitry, which requires $O(\sqrt[3]{N})$ control lines for $N$ frequency-tunable qubits Bunyk2014 .
(ST1) This control scheme comprises a microwave drive pulse and a baseband flux pulse. Similar to Refs. Bejanin2022 ; Zhao2023 , single-qubit gates, such as $\sqrt{X}$ gates, are realized by individually tuning qubits on resonance with the shared microwave drive. However, unlike the previous works using an always-on microwave drive, here a microwave drive pulse is employed (compared to that with the always-on drive Zhao2023 , this shared control pulse will not affect other qubit operations, such as qubit readout, qubit reset, and two-qubit gates). Thus the derivative removal by adiabatic gate (DRAG) can be used for realizing single-qubit gates Motzoi2009 . Accordingly, Figure 4(a) shows the typical two-pulse configuration, which consists of a raised cosine flat-top flux pulse for tuning the qubit frequency from the idle point to the working point and a cosine DRAG pulse for XY control Chen2016 . As qubit idling frequencies should be different from qubits to qubits, the amplitude of the flux pulse applied to each qubit should be individually tuned on resonance with the shared microwave drive. Given that the flux pulse is applied globally to qubits through the shared Z lines, individual tuning can be achieved by realizing tunable coupling between qubits and the shared Z lines. Following Ref. Johnson2010 , Figure 3(a) shows the tunable rf-SQUID coupler Harris2009 ; Brink2005 ; Harris2007 , which itself is controlled by a $\Phi$-DAC, for individually tuning the amplitudes of flux pulses felt by qubits.
(ST2) This scheme consists of a parametric drive pulse through the Z line and a microwave drive pulse through the XY line. As illustrated in Refs.Li2021 ; Caldwell2018 ; Strand2013 , modulating the qubit frequency with a parametric drive can induce a series of sidebands due to the nonlinear dependence of the qubit frequency on the flux bias. When one of such sidebands is on-resonance with the microwave drive, coherent qubit control can be achieved Li2021 . Figure 4(b) shows such a pulse configuration, in which a flat-top parametric drive pulse with cosine-shape ramps is used to modulate the qubit frequency and a raised cosine flat-top microwave drive pulse is for XY control. Note that similar schemes have also been studied recently in semiconductor spin qubits Gyorgy2022 ; Bosco2023 ; John2023 .
Type-SF: This category is for single-qubit gates applied to fixed-frequency qubits. During single-qubit gates, a pair of microwave drive pulses are simultaneously delivered to the qubit through the two XY lines, as shown in Fig. 3(b).
(SF1) Based on an on-chip ’three-wave mixer’ Frattini2017 ; Chapman2023 ; Liu2014 ; Zhao2018 , the two microwave drives are converted to an effective qubit drive. When the sum-frequency or difference-frequency of the two drives is on-resonance with the qubit, coherent control of the qubit can be achieved with the effective drive. Figure 3(b) schematically illustrates such an on-chip mixer based on flux qubit Liu2014 ; Zhao2018 , which is capacitively coupled to the transmon qubit and is driven by the two microwave drives. Here, for example, we consider that the difference frequency of pulses equals the qubit frequency. Accordingly, Figure 4(c) shows the typical used raised cosine flat-top pulses.
(SF2) Here, single-qubit gates are realized by applying two off-resonance microwave drives to the qubit Dai2017 ; He2019 ; Koong2021 ; Bracht2021 ; Wang2023 , as shown in Fig. 3(b). Such scheme can be understood as follows: one of the two off-resonance drives is used to dress the (bare) qubit and shift the qubit frequency through the ac-Stark effect, while the second one is applied for controlling such microwave-dressed qubit Zhao2022c ; Bracht2023 ; Zuk2023 . Moreover, this two-tone drive scheme can be combined with the DRAG scheme to suppress leakage errors during gate operations. Figure 4(d) shows the typical used cosine DRAG pulses.
@@
Figure 5: The numerical verification of the four single-qubit gate schemes. In the numerical analysis, the transmon qubit is treated as an anharmonic oscillator Koch2007 with the (maximum) frequency of $\omega_{q}/2\pi=5.25\,\rm GHz$ and the anharmonicity of $\eta_{q}/2\pi=-250\,\rm MHz$. The upper panel shows the population $P_{0}$ at the end of the applied gate pulses shown in Fig. 4 with the qubit prepared in state $|0\rangle$, while the bottom panel is for the leakage $L_{1}$ (white regions are where leakage below $10^{-5}$). Dashed grey lines indicate the available regions for realizing $\sqrt{X}$ gates while the red star is for $X$ gates. (a) $P_{0}$ and $L_{1}$ versus the peak drive amplitude $\Omega$ of the DRAG pulse (with the frequency of $\omega_{d}/2\pi=5.10\,\rm GHz$) and the frequency bias $\Delta_{b}=\omega_{q}-\omega_{d}$. (b) $P_{0}$ and $L_{1}$ versus the parametric drive frequency $\omega_{p}$ and the microwave drive amplitude $\Omega$ with the drive frequency of $\omega_{d}/2\pi=5.75\,\rm GHz$. Here, the dependence of qubit frequency on the flux bias is approximated by $\omega(\Phi)=(\omega_{q}-\eta_{q})\sqrt{|\cos(\pi\Phi/\Phi_{0})|}+\eta_{q}$ Koch2007 , where $\Phi_{0}$ denotes the flux quantum. The used parametric drive is $\Phi=0.05\Phi_{0}+\Phi_{p}(t)\Phi_{0}\cos(\omega_{p}t+\phi_{0})$ (for simplicity, here $\phi_{0}=0$) with the peak drive amplitude $\Phi_{p}=0.2$. The result shows the first-order sideband transition with $\omega_{p}\sim|\omega_{d}-\omega_{q}|$. (c) $P_{0}$ and $L_{1}$ versus the peak drive amplitudes $\Omega$ of the two pulses and the frequency of the second drive $\omega_{2}$. The frequency of the first drive is $\omega_{1}/2\pi=3.5\,\rm GHz$. Here, the flux qubit which is treated as a qutrit with the cycle transitions of $(\omega_{01}/2\pi=6.25\,{\rm GHz},\,\omega_{02}/2\pi=10\,{\rm GHz},\,\omega_{12}/2\pi=3.75\,\rm GHz)$ and the transmon-flux coupling strength is $(g_{01}/2\pi=94\,{\rm MHz},\,g_{02}/2\pi=140\,{\rm MHz},\,g_{12}/2\pi=136\,{\rm MHz})$  Zhao2018 . (d) $P_{0}$ and $L_{1}$ versus the drive amplitude ratio $\Omega_{2}/\Omega_{1}$ and the drive-qubit detuning ration $\delta_{2}/\delta_{1}$ of the two pulses with $\delta_{1(2)}=\omega_{1(2)}-\omega_{q}$. Here, for the first drive, the peak drive parameters and the drive-qubit detuning are $\Omega_{1}/2\pi=15\rm MHz$ and $\delta_{1}/2\pi=50\,\rm MHz$, respectively.
For each of the schemes schematically illustrated in Fig. 3 (also summarized in Table 1), Figure 5 presents the numerical verification of their feasibility for performing single-qubit gates, e.g., $\sqrt{X}$ gates or $X$ gates, according to the control pulses shown in Fig. 4. Here, the upper panel shows the population ($P_{0}$) in state $|0\rangle$ at the end of the applied pulses with the qubit prepared in its ground state. The dashed grey line denotes the available regions of the control parameters for realizing $\sqrt{X}$ gates while the red star is for $X$ gates. Furthermore, the bottom panel shows the leakage $L_{1}$ Wood2018 . These numerical results suggest that low-leakage, high-fidelity single-qubit X rotations can be achieved with the proposed schemes.
As shown in Fig. 2(a), to achieve universal single-qubit control, here we consider compiling arbitrary single-qubit gates into two $\sqrt{X}$ gates and three physical (or virtual) $Z$ gates McKay2017 . The main reason for taking such a compiling method is twofold. First, we prefer the $\sqrt{X}$ gate rather than the $X$ gate due to that the former one can provide more flexibility in choosing control parameters. This could provide possible adaptations to inhomogeneities in coupling efficiencies between qubits and control lines, defects affecting the qubit performance Zhao2023 , frequency crowding in choosing control frequencies, and so on. Second, as control pulses are shared by qubits, individual Z gate applied to each qubit cannot be directly realized by the phase shift of the shared microwave drive or parametric drive McKay2017 ; Chen2023 , especially in the gate schemes of (ST1), (ST2), and (SF2). Besides the main reason, single-qubit errors, that result from microwave crosstalk, could be mitigated by optimizing the Z gates Wang2022 ; Wang2022a .
@@
Figure 6: Single-qubit Z gates. The arbitrary Z gates can be obtained by applying two off-resonance drives. Here, similar to Fig. 4(c), two 150-ns flat-top pulses with a ramp time of 50 ns are utilized. The left panel shows the angles of the Z rotation versus the two drive-qubit detunings $\delta_{1}$ and $\delta_{2}$. The associated Z gate error is shown in the right panel (white regions are where gate errors are below $10^{-4}$).
@@
For the gate scheme of (SF1) based on the tree-wave mixing process, the phase of the effective drive applied to the qubit can be controlled by the difference phase or sum phase of the two pulses. Thus, in principle, the virtual Z gates can be utilized here. However, as mentioned above, for the gate schemes of (ST1), (ST2), and (SF2), the physical Z gates are needed. Generally, this can be realized by using the ac-Stark effect resulting from off-resonance microwave drives Wei2023 . Here, we consider the implementation of arbitrary Z gates in (SF2), where two off-resonance microwave drive pulses are applied to the qubit. Similar to Fig. 4(c), a pair of raised cosine flat-top pulses are employed for performing Z rotations. Figures 6(a) and 6(b) show the Z rotation angle and the gate fidelity Pedersen2007 versus the two drive-qubit detunings, respectively, illustrating that arbitrary Z rotations can be realized with high fidelity. Note here that while similar to the gate scheme of (F1), the phase difference can be controlled, the nonlinear dependence of the qubit dynamics on the frequency difference or phase difference makes it incompatible with the virtual Z scheme Dai2017 , see Appendix A for details.

### 20 Mar 2024 v2

@@ III Gate operations in the multiplexed control architecture
Here we turn to illustrate how single- and two-qubit gates are realized in the multiplexed control architecture. As shown in Figs. 1(c) and 2, gate operations, unlike that in the conventional independent control scheme, are implemented by applying simultaneous control pulses to all shared row and column lines. When focusing on one particular qubit or coupler within the lattice, single- and two-qubit gate operations are realized by applying a unique pair of control pulses to a qubit or a coupler. Within this context, we propose various gate schemes that are both compatible with the two-pulse configuration and the existing superconducting qubit technologies. For illustration purposes, we focus on the gates applied to transmon qubits Koch2007 , but in principle, it should also be feasible for other superconducting qubits, such as fluxonium qubits Manucharyan2009 .
@@ III.1 Single-qubit gates
Table 1: Single-qubit gate schemes with two control pulses. The schematic illustration of these schemes is shown in Fig. 3. Note that for frequency-tunable qubits, static dc flux biases (not listed here but shown in Fig. 3) are generally required to bias the qubits at their idle points.
@@
Figure 5: The numerical verification of the four single-qubit gate schemes. In the numerical analysis, the transmon qubit is treated as an anharmonic oscillator Koch2007 with the (maximum) frequency of $\omega_{q}/2\pi=5.25\,\rm GHz$ and the anharmonicity of $\eta_{q}/2\pi=-250\,\rm MHz$. The upper panel shows the population $P_{0}$ at the end of the applied gate pulses depicted in Fig. 4, with the qubit prepared in state $|0\rangle$, while the bottom panel illustrates the leakage $L_{1}$ (white regions indicate leakage below $10^{-5}$). Dashed grey lines indicate the available regions for realizing $\sqrt{X}$ gates while the red star is for $X$ gates. (a) $P_{0}$ and $L_{1}$ versus the peak drive amplitude $\Omega$ of the DRAG pulse (with a frequency of $\omega_{d}/2\pi=5.10\,\rm GHz$) and the frequency bias $\Delta_{b}=\omega_{q}-\omega_{d}$. (b) $P_{0}$ and $L_{1}$ versus the parametric drive frequency $\omega_{p}$ and the microwave drive amplitude $\Omega$ with the drive frequency of $\omega_{d}/2\pi=5.75\,\rm GHz$. Here, the dependence of qubit frequency on the flux bias is approximated by $\omega(\Phi)=(\omega_{q}-\eta_{q})\sqrt{|\cos(\pi\Phi/\Phi_{0})|}+\eta_{q}$ Koch2007 , where $\Phi_{0}$ denotes the flux quantum. The used parametric drive is $\Phi=0.05\Phi_{0}+\Phi_{p}(t)\Phi_{0}\cos(\omega_{p}t+\phi_{0})$ (for simplicity, here $\phi_{0}=0$) with a peak drive amplitude of $\Phi_{p}=0.2$. The result shows the first-order sideband transition with $\omega_{p}\sim|\omega_{d}-\omega_{q}|$. (c) $P_{0}$ and $L_{1}$ versus the peak drive amplitudes $\Omega$ of the two pulses and the frequency of the second drive $\omega_{2}$. The frequency of the first drive is $\omega_{1}/2\pi=3.5\,\rm GHz$. Here, the flux qubit is modeled as a qutrit with cycle transitions $(\omega_{01}/2\pi=6.25\,{\rm GHz},\,\omega_{02}/2\pi=10\,{\rm GHz},\,\omega_{12}/2\pi=3.75\,\rm GHz)$ and transmon-flux coupling strength $(g_{01}/2\pi=94\,{\rm MHz},\,g_{02}/2\pi=140\,{\rm MHz},\,g_{12}/2\pi=136\,{\rm MHz})$  Zhao2018 . (d) $P_{0}$ and $L_{1}$ versus the drive amplitude ratio $\Omega_{2}/\Omega_{1}$ and the drive-qubit detuning ration $\delta_{2}/\delta_{1}$ of the two pulses, where $\delta_{1(2)}=\omega_{1(2)}-\omega_{q}$. Here, for the first drive, the peak drive parameters and the drive-qubit detuning are $\Omega_{1}/2\pi=15\rm MHz$ and $\delta_{1}/2\pi=50\,\rm MHz$, respectively.
@@
Figure 6: Single-qubit Z gates. Z gates can be obtained by utilizing two off-resonance drives. Similar to Fig. 4(c), two 150-ns flat-top pulses with a ramp time of 50 ns are utilized. The left panel shows the angles of the Z rotation versus the two drive-qubit detunings $\delta_{1}$ and $\delta_{2}$. Meanwhile, the associated Z gate error is shown in the right panel, with white regions indicating gate errors below $10^{-4}$.
Table 1 summarizes four single-qubit gate schemes using two control pulses, which are also schematically illustrated in Fig. 3. These schemes can be grouped into two categories, depending on whether the qubit is tunable. In the following, we will delve into detailed explanations of each of the four gate schemes.
Type-ST: This category pertains to single-qubit gates designed for frequency-tunable qubits. During gate operations, a flux pulse and a microwave pulse are simultaneously delivered to the qubit through a shared Z line and a shared XY line, respectively, as shown in Fig. 3(a). Furthermore, for frequency-tunable qubits, static dc flux biases are generally needed for biasing the qubits at their idle points. Following Ref. Johnson2010 , this flux offset can in principle be introduced by using a $\Phi$-DAC, which functions as on-chip programmable magnetic memory and provides the required static bias. Moreover, the $\Phi$-DAC can be efficiently programmed and addressed by the addressing circuitry, which requires $O(\sqrt[3]{N})$ control lines for $N$ frequency-tunable qubits Bunyk2014 .
(ST1) This control scheme involves a microwave drive pulse and a baseband flux pulse. Similar to Refs. Bejanin2022 ; Zhao2023 , single-qubit gates, such as $\sqrt{X}$ gates, are realized by individually tuning qubits on resonance with the shared microwave drive. Unlike previous works using an always-on microwave drive Zhao2023 , here a microwave drive pulse is used to prevent any negative effects on other qubit operations, such as qubit readout, qubit reset, and two-qubit gates. Thus the derivative removal by adiabatic gate (DRAG) can be employed for realizing single-qubit gates Motzoi2009 . Figure 4(a) shows the typical two-pulse configuration, which consists of a raised cosine flat-top flux pulse for tuning the qubit frequency from the idle point to the working point and a cosine DRAG pulse for XY control Chen2016 . As qubit idling frequencies should be different from qubits to qubits, the amplitude of the flux pulse applied to each qubit should be individually fine-tuned to resonate with the shared microwave drive. Given that the flux pulse is applied globally to qubits through the shared Z lines, individual tuning is required and can be achieved by realizing tunable coupling between qubits and the shared Z lines. Following Ref. Johnson2010 , Figure 3(a) shows the tunable rf-SQUID coupler Harris2009 ; Brink2005 ; Harris2007 , which itself is controlled by a $\Phi$-DAC, for individually tuning the amplitudes of flux pulses felt by qubits.
(ST2) This scheme entails a parametric drive pulse through the Z line and a microwave drive pulse through the XY line. As illustrated in Refs.Li2021 ; Caldwell2018 ; Strand2013 , modulating the qubit frequency with a parametric drive can induce a series of sidebands due to the nonlinear dependence of the qubit frequency on the flux bias. When one of these sidebands is on-resonance with the microwave drive, it enables coherent qubit control Li2021 . Figure 4(b) shows this pulse configuration, where a flat-top parametric drive pulse with cosine-shape ramps modulates the qubit frequency and a raised cosine flat-top microwave drive pulse facilitates XY control. Note that similar schemes have also been studied recently in semiconductor spin qubits Gyorgy2022 ; Bosco2023 ; John2023 .
Type-SF: This category pertains to single-qubit gates intended for fixed-frequency qubits. During single-qubit gates, a pair of microwave drive pulses are simultaneously delivered to the qubit through the two XY lines, as shown in Fig. 3(b).
(SF1) Leveraging an on-chip ’three-wave mixer’ Frattini2017 ; Chapman2023 ; Liu2014 ; Zhao2018 , the two microwave drives are converted to an effective qubit drive. When the sum-frequency or difference-frequency of the two drives is on-resonance with the qubit, coherent control of the qubit can be achieved with the effective drive. Figure 3(b) schematically illustrates such an on-chip mixer based on flux qubit Liu2014 ; Zhao2018 , which is capacitively coupled to the transmon qubit and is driven by the two microwave drives. Here, for example, we consider that the difference frequency of pulses equals the qubit frequency. Accordingly, Figure 4(c) shows the typical used raised cosine flat-top pulses.
(SF2) In this scenario, single-qubit gates are realized by applying two off-resonance microwave drives to the qubit Dai2017 ; He2019 ; Koong2021 ; Bracht2021 ; Wang2023 , as shown in Fig. 3(b). Such scheme can be understood as follows: one of the two off-resonance drives is used to dress the (bare) qubit and shift the qubit frequency through the ac-Stark effect, while the second one is applied for controlling such a microwave-dressed qubit Zhao2022c ; Bracht2023 ; Zuk2023 . Moreover, this two-tone drive scheme can be combined with the DRAG scheme to suppress leakage errors during gate operations. Figure 4(d) depicts the typical used cosine DRAG pulses in this context.
For each of the schemes depicted in Fig. 3 (also summarized in Table 1), Figure 5 presents the numerical verification of their feasibility for performing single-qubit gates, e.g., $\sqrt{X}$ gates or $X$ gates, according to the control pulses shown in Fig. 4. Here, the upper panel shows the population ($P_{0}$) in state $|0\rangle$ at the end of the applied pulses with the qubit prepared in its ground state. The dashed grey line denotes the feasible regions of the control parameters for realizing $\sqrt{X}$ gates while the red star is for $X$ gates. Furthermore, the bottom panel shows the leakage $L_{1}$ Wood2018 . These numerical findings indicate that that low-leakage, high-fidelity single-qubit X rotations can be achieved with the proposed schemes.
As shown in Fig. 2(a), to achieve universal single-qubit control, we consider compiling arbitrary single-qubit gates into two $\sqrt{X}$ gates and three physical (or virtual) $Z$ gates McKay2017 . The main reason for taking this compiling method is twofold. First, we prefer the $\sqrt{X}$ gate rather than the $X$ gate due to that the former one can provide more flexibility in choosing control parameters. This flexibility could provide potential adaptations to address challenges involving inhomogeneities in coupling efficiencies between qubits and control lines, defects affecting qubit performance Zhao2023 , frequency crowding in choosing control frequencies, and so on. Second, as control pulses are shared by qubits, individual Z gate applied to each qubit cannot be directly realized by the phase shift of the shared microwave drive or parametric drive McKay2017 ; Chen2023 , especially in the gate schemes of (ST1), (ST2), and (SF2). Besides the two main reason, single-qubit errors, that result from microwave crosstalk, could be mitigated by optimizing the Z gates Wang2022 ; Wang2022a in this decomposition.
Figure 7: The numerical verification of the two-qubit gate operations based on two-tone flux modulations. In the numerical analysis, we consider that two transmon qubits, labeled by $Q_{a}$ and $Q_{b}$, are coupled via a tunable bus, labeled by $Q_{bus}$, and the system state is denoted by $|Q_{a}Q_{bus}Q_{b}\rangle$. The qubit frequencies are $\omega_{a(b)}/2\pi=5.05(5.25)\,\rm GHz$ and the qubit anharmonicity is $\eta_{a(b)}/2\pi=-250\,\rm MHz$. The maximum frequency and the anharmonicity of the tunable bus are $\omega_{bus}/2\pi=6.20\,\rm GHz$ and $\eta_{bus}/2\pi=-200\,\rm MHz$, respectively. The qubit-bus coupling strength is $g/2\pi=100\,\rm MHz$. The dependence of bus frequency on the two-tone flux modulation is approximated by $\omega_{bus}(\Phi)=(\omega_{bus}-\eta_{bus})\sqrt{|\cos(\pi\Phi(t)/\Phi_{0})|}+\eta_{bus}$ with $\Phi(t)=\Phi_{p}(t)[\sin(\omega_{p1}t)+\sin(\omega_{p2}t+\phi_{0})]$ (where $\phi_{0}=0$ for simplicity) and the modulation frequency $\omega_{p1}/2\pi=2.0\,\rm GHz$. Here, similar to Fig. 4(c), two raised cosine flat-top pulses are used and the peak modulation amplitude is $\Phi_{p}$. (a) Population $P_{100}$ (upper panel) and $P_{001}$ (bottom panel) versus the modulation amplitude $\Phi_{p}$ and the modulation frequency $\omega_{p2}$ with the qubit prepared in state $|100\rangle$. The ramp time is 15 ns and the total gate length is 100 ns. Dashed orange lines indicate the available regions for realizing $\sqrt{\textmd{iSWAP}}$ gates while the red star is for iSWAP gates. Here, the two-qubit gate is actuated when the difference frequency of the two modulations matches the qubit-qubit detuning. (b) Population $P_{101}$ (upper panel) versus the modulation amplitude $\Phi_{p}$ and the modulation frequency $\omega_{p2}$ with the system prepared in state $|101\rangle$. The bottom panel shows the associated conditional phase. The ramp time is 25 ns and the total gate length is 200 ns. Here, the CZ gate is realized when the difference frequency of the two modulations matches the detuning between the energy levels of $|101\rangle$ and $|200\rangle$. Note that discontinuities near the horizontal line at $\Phi_{p}/\Phi_{0}=0.13$ arise from the parasitic interaction between the energy levels of $|101\rangle$ and $|110\rangle$.
@@
For the gate scheme of (SF1) based on the tree-wave mixing process, the phase of the effective drive applied to the qubit can be controlled by the difference phase or sum phase of the two pulses. Thus, in principle, the virtual Z gates can be utilized here. However, as mentioned above, the gate schemes of (ST1), (ST2), and (SF2) necessitate physical Z gates. Generally, this can be realized by using the ac-Stark effect due to off-resonance drives Wei2023 . Here, we consider the implementation of arbitrary Z gates within (SF2), where two off-resonance microwave drive pulses are applied to the qubit. Similar to Fig. 4(c), a pair of raised cosine flat-top pulses are employed for performing Z rotations. Figures 6(a) and 6(b) show the Z rotation angle and the gate fidelity Pedersen2007 versus the two drive-qubit detunings, respectively, illustrating that arbitrary Z rotations can be realized with high fidelity. Note here that while similar to the gate scheme of (F1) the phase difference can be controlled, the nonlinear dependence of the qubit dynamics on the frequency difference or phase difference makes it incompatible with the virtual Z scheme Dai2017 , see Appendix A for details.